OpenStudy (anonymous):

In a geometric series, if the sum of the first 3 terms is -3/8 and the sum to infinity is -1/3, what is the common ratio?

OpenStudy (bradely):

sn =a(1-r^n)/(1-r) =sinf(1-r^n) since sinf =a/(1-r) (-3/8)=(-1/3)(1-r^3) (1-r^3)=9/8 r^3=-1/8 r=-1/2 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

\[S_{\infty}=\frac{1}{1-r}\]

OpenStudy (anonymous):

plug in what you know \(S_{\infty}=-\large{\frac{1}{3}}\) and solve for r (the common ratio)

OpenStudy (anonymous):

oops, should be \[S_\infty=\frac{ a_n }{ 1-r }\]

OpenStudy (anonymous):

can use this and the fact that \(S_3 = -\large{\frac{3}{8}}\) to find r

OpenStudy (anonymous):

Let the first term be a and common ratio r. \[a+ar+ar^2=-\frac{ 3 }{ 8 },a \left( 1+r+r^2 \right)=-\frac{ 3 }{ 8 }\] \[S _{\infty}=\frac{ a }{ 1-r },-\frac{ 1 }{ 3 }=\frac{ a }{ 1-r },a=-\frac{ 1-r }{ 3 }\] \[-\frac{ 1-r }{ 3 }\left( 1+r+r^2 \right)=-\frac{ 3 }{ 8 }\] \[1-r^3=-\frac{ 3 }{ 8 }\times- 3=\frac{ 9 }{ 8 },-r^3=\frac{ 9 }{ 8 }-1=\frac{ 9-8 }{ 8 }=\frac{ 1 }{ 8 }\] \[r^3=-\frac{ 1 }{ 8 }=\left( \frac{ -1 }{ 2 } \right)^3,r=-\frac{ 1 }{ 2 }\]